OpenStudy (anonymous):

integral of (cosX+sin2X)/sinX

OpenStudy (turingtest):

\[\int\frac{\cos x+\sin(2x)}{\sin x}dx\]?

OpenStudy (turingtest):

\[\sin(2x)=2\sin x\cos x\]then simplify the fraction. Shouldn't be too hard to integrate after that.

OpenStudy (turingtest):

are you stuck?

OpenStudy (anonymous):

Thanks! i guess i just figured it out!

OpenStudy (turingtest):

cool :)

OpenStudy (anonymous):

I have another one... : l

OpenStudy (turingtest):

could you post it separately please?